OpenStudy (meilendurcer):

Choose the slope-intercept equation of the line that passes through the point (4, -6) and is parallel to y = -one halfx + 5. y = -2x + 2 y = 2x - 14 y = one halfx - 8 y = -one halfx - 4

OpenStudy (anonymous):

For two lines to be parellel both should have same slope nad substitute point in options, u will get solution

OpenStudy (anonymous):

slope is coefficient of x

OpenStudy (meilendurcer):

i wanna say the last one but i'm not sure if that right

ganeshie8 (ganeshie8):

yes so start by figuring out the slope of given line : \( y = \frac{1}{2}x +5\)

ganeshie8 (ganeshie8):

thats a very good guess, but no

OpenStudy (anonymous):

what @meilendurcer said is right slope is -1/2

ganeshie8 (ganeshie8):

we can tell the slope of given line just by looking at its equation : \(y = \color{red}{m}x+b\) slope = \(\color{red}{m}\)

OpenStudy (meilendurcer):

y = -\[y= -\frac{ 1 }{ 2 } +5\]

OpenStudy (meilendurcer):

forgot to delte the top part

ganeshie8 (ganeshie8):

its okay :) thats the given line, so whats the slope ?

ganeshie8 (ganeshie8):

\(\large y = \color{red}{-\frac{1}{2}}x+5\) slope ?

OpenStudy (meilendurcer):

\[y= - \frac{ 1 }{ 2 } x -4\]

OpenStudy (meilendurcer):

yes -1/2 is the slope

OpenStudy (anonymous):

Yes thats the correct answer

ganeshie8 (ganeshie8):

cool :) thats the right answer indeed ! was just checking if you get it properly ;p

OpenStudy (meilendurcer):

@ganeshie8 yay! please help with another??

ganeshie8 (ganeshie8):

sure ask

OpenStudy (meilendurcer):

did we do this one? The functions f(x) and g(x) are described below: f(x) = 2x + 9 g(x) = 2x - 15 The graph of g(x) is obtained by shifting down the graph of f(x) by _____ units.

ganeshie8 (ganeshie8):

any good guesses this time ? :)

OpenStudy (meilendurcer):

no not yet

OpenStudy (meilendurcer):

unless

OpenStudy (meilendurcer):

f(x) = 2x15 +9 g(x) = 2x9 - 15?

ganeshie8 (ganeshie8):

lets try and write g(x) in terms of f(x) : g(x) = 2x - 15 = 2x + 9 -9 - 15 = 2x + 9 - 24

ganeshie8 (ganeshie8):

g(x) = f(x) - 24

ganeshie8 (ganeshie8):

Clearly g(x) is a 24 units shifted down version of f(x) ^

OpenStudy (meilendurcer):

ohh, well that was a wild guess. I still need practice to really learn those.

OpenStudy (meilendurcer):

thank you!

ganeshie8 (ganeshie8):

hahah we can tame the wilderness by practice ;) yw !

OpenStudy (meilendurcer):

@ganeshie8 lol yeah

OpenStudy (skullpatrol):